Specify the Electric Field Strengths at different distances?

A 10-cm-long thin glass rod uniformly charged to 5.00 nC and a 10-cm-long thin plastic rod uniformly charged to - 5.00 nC are placed side by side, 4.00 cm apart. Whatare the electric field strengths to at distances 1.0 cm, 2.0 cm, and 3.0 cm from the glass rod along the line connecting the midpoints of the two rods?

E1:
E2:
E3:
Answer #1

q = 5.00 nC, L = 10 cm, d = 4.00 cm, r = 1.0 cm, or 2.0 cm, or 3.0 cm

E(r) = q/(2πε0rL) + q/[2πε0(d - r)L] = q/(2πε0L) * [1/r + 1/(d - r)] = 900 [1/r + 1/(d - r)]

E1=E(1.0 cm) = 1.20*105N/C

E2=E(2.0 cm) = 9.00*104N/C

E3=E(3.0 cm) = 1.20*105N/C

answered by: Momito